Topic: Normal Distribution Distribution & Inverse Normal Distribution


1.) The continuous random variable Z has a standard normal distribution, N (0,1). Find the following probabilities by using tables.

a.) P(Z>2)

b.) P(Z<1.35)

c.) P(-0.8<Z<1.8)



1
Expert's answer
2021-07-16T13:06:49-0400

a) "P(Z>2)=1-P(Z<2)=1-0.9772=0.0228."

b) "P(Z<1.35)=0.9115."

c) "P(-0.8<Z<1.8)=P(Z<1.8)-P(Z<-0.8)="

"=0.9641-0.2119=" 0.7522.